OpenStudy (anonymous):

ODE xy'+(2x-3)y=4x^4 After I get to e^2x x^-3 I get lost.

sam (.sam.):

separable xy'+(2x-3)=4x^4 \[y'=\frac{4x^4-2x+3}{x} \\ \\ \int\limits \frac{4x^4-2x+3}{x} dx\] \[x^4-2x+3\ln(x)+c\]

OpenStudy (anonymous):

I wrote it wrong, it's xy'+(2x-3)y=4x^4

OpenStudy (sirm3d):

after finding the integrating factor, which is \[e^{2x} x^{-3}\] you now solve the equation \[(e^{2x}x^{-3})y=\int(e^{2x}x^{-3})(4x^4)dx\]

OpenStudy (sirm3d):

in the ODE \[y'+P(x)y=Q(x)\] the integrating factor is \[v=e^{\int P(x) dx}\] and the solution to the ODE is given by \[vy=\int v\cdot Q(x) dx\]

OpenStudy (anonymous):

Thanks!